OpenStudy (anonymous):

What are the asymptotes of the hyperbola given by the equation y squared over 1minus x squared over 121 = 1?

OpenStudy (rajee_sam):

\[\frac{ y ^{2} }{ 1 } - \frac{ x ^{2} }{ 121 } = 1\] This is of the form \[\frac{ y ^{2} }{ a ^{2} } - \frac{ x ^{2} }{ b ^{2} } = 1\] This Hyperbola will have a graph like the following with ots vertex in the y axis.

OpenStudy (rajee_sam):

Now \[a ^{2} = 1; \space so \space a = 1\]\[b ^{2} = 121; \space So \space b = 11\] The asymptotes are given by the equation \[y = \pm \frac{ a }{ b } x\] \[y = \pm \frac{ 1 }{ 11 }x\]